OpenStudy (anonymous):

Kelly’s scores on her first two tests in Algebra were 52 and 89. What will she need to make on third test to have an average of at least 79? Answer 96 or better 85 or better 79 or better 80 or better

OpenStudy (queelius):

The average is defined as the sum of the scores divided by the number of scores, e.g., if you get a 100 on test 1, 80 on test 2, and 90 on test 3, your average is: (100 + 80 + 90) / 3 = 270 / 3 = 90 So, we want to solve the following problem: \[\frac{ 52 + 89 + x }{ 3 } \ge 79\]

OpenStudy (queelius):

Solve for x.